OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3). A. 2 B. 5 C. 8 D. -2

OpenStudy (anonymous):

x^2 or 2x

OpenStudy (anonymous):

what is mean by x2 x^2 or 2x

OpenStudy (anonymous):

Oh, sorry. I meant x^2...

OpenStudy (anonymous):

C

OpenStudy (anonymous):

\[f(x)=x^2-1\] \[f(3)=3^2-1\] =9-1 =8